Sherds of clay vessels were put together to reconstruct rim diameters of the original ceramic vessels at the Wind Mountain archaeological site†. A random sample of ceramic vessels gave the following rim diameters (in centimeters).

15.9 13.4 22.1 12.7 13.1 19.6 11.7 13.5 17.7 18.1

(a) Use a calculator with mean and sample standard deviation keys to find the sample mean x and sample standard deviation s. (Round your answers to one decimal place.)

x = cm
s = cm


(b) Compute a 95% confidence interval for the population mean μ of rim diameters for such ceramic vessels found at the Wind Mountain archaeological site. (Round your answers to one decimal place.)

lower limit     cm
upper limit     cm
Answer #1

Solution:

x x2
15.9 252.81
13.4 179.56
22.1 488.41
12.7 161.29
13.1 171.61
19.6 384.16
11.7 136.89
13.5 182.25
17.7 313.29
18.1 327.61
\sumx=157.8 \sumx2=2597.88

a ) The sample mean is \bar x

Mean \bar x     = (\sumx / n) )

= (15.9+13.4+22.1+12.7+13.1+19.6+11.7+13.5+17.7+18.1/ 10 )

=157.8 / 14

= 15.78

Mean \bar x    = 15.78

The sample standard is S

  S = \sqrt{} ( \sum x2 ) - (( \sum x)2 / n ) n -1

= \sqrt{} (2597.88( (157.8)2 / 10 ) 9

   = \sqrt{} ( 2597.88- 2490.084 / 9)

=\sqrt{} (107.796/ 9)

= \sqrt{} 11.9773

= 3.4608

The sample standard is = 3.46

b ) Degrees of freedom = df = n - 1 = 10- 1 = 9

At 95% confidence level the t is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,9 = 2.262

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.262* (3.46 / \sqrt 10)

= 2.47

Margin of error = 2.47

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

15.78 - 2.47 < \mu < 15.78 + 2.47

13.31 < \mu < 18.25

(13.31, 18.25 )